Random samples of size 3 are taken from a population of the numbers 3, 4, 5, 6, 7, 8, and 9.

1. How many samples are possible? 



2. Construct the sampling distribution of the sample means. 


1
Expert's answer
2022-04-12T13:17:55-0400
  1. Consider all samples with replacement. We get: "7^3=343" possible samples. For samples without replacement we get: "7\cdot6\cdot5=210."
  2. We consider sampling distribution of the sample means without replacements. We receive samples of the type: "(i,j,k)," where "i\neq j\neq k". The probability of the mean "\frac{i+j+k}{3}" is: "\frac{1}{210}\cdot N", where "N" is a number of samples with the same mean. The number can be found either from the table with samples or from equality "i_1+j_1+k_1=i_2+j_2+k_2" that has to be considered for all samples.